Finch Computer Services, Inc. has been in business for six months. The following are basic operating data for that period Month July 114 Aug. Sept Oct. Nov. Dec. Service hours Revenue Operating costs 328 $6,270 $7,370 $14,190 $23,540 $17,600 $18,040 $4,260 $5,280 7,170 $11,170 9,150 $10,550 134 258 428 320 Required a. What is the average service revenue per hour in each month and the overall average for the six-month period? b. Use the high-low method to estimate the total monthly fixed cost and the variable cost per hour c. Determine the average contribution margin per hour. Complete this question by entering your answers in the tabs below. Required A Required B Required C What is the average service revenue per hour in each month and the overall average for the six-month period? Month July Aug Sept. Oct. Nov Dec. Total Average Average service revenue per hourMonth July 114 Sept. 258 Aug. Oct. Nov. Dec. Service hours Revenue Operating costs 328 $6,270 $7,370 $14,190 $23,540 $17,600 $18,040 $4,260 $5,280 7,170 $11,170 9,150 $10,550 134 428 320 Required a. What is the average service revenue per hour in each month and the overall average for the six-month period? b. Use the high-low method to estimate the total monthly fixed cost and the variable cost per hour. c. Determine the average contribution margin per hour. Complete this question by entering your answers in the tabs below Required A Required B Required C Use the high-low method to estimate the total monthly fixed cost and the variable cost per hour. (Do not round intermediate calculations. Round Variable cost to 2 decimal places.) Variable cost per hour Fixed cost per month
Answer #1

A. Calculation of average revenue per hour :

Average revenue per hour = Revenue/ Service Hours

July August September October November December Total
Revenue 6270 7370 14190 23540 17600 18040 87010
Service Hours 114 134 258 428 320 328 1582
Average service revenue per hour 55 55 55 55 55 55 55

B. Using High Low Method Monthly Fix Cost & Variable Cost calculation :-

Month Operating Cost Service Hours Variable Cost
Oct high 11170 428
July Low 4260 114
Difference 6910 314 (6910/314)        22.00

Fixed Cost = Operating Cost in October - (Variable cost per unit * Service   hours in October)

= 11170 - (428*22.00)

= 11170-9416 = 1754

Therefore, Monthly Fix Cost = $1754  

Variable Cost Per Hour = $22.00

C. Average Contribution margin per hour

Particulars Amount
Average Revenue per hour(As Calculate in 55
(Less)Variable cost per hour 22
Average Contribution margin per hour 33